OpenStudy (anonymous):

Which of the following represents x =1/3y written in general form? 1/3x - y = 0 3x - y = 0 x - 1/3y = 0

OpenStudy (zepp):

The general form is when one side is set as 0 \[x = \frac{1}{3}y\]\[x-\frac{1}{3}y = \frac{1}{3}y-\frac{1}{3}y\]\[x-\frac{1}{3}y=0\]